2.5 Teresa spends her entire budget and consumes 6 units of z and 20 units of y. The price ofr is twice the price of y. Her income doubles and the price of y doubles, but the price of z stays the same. If she continues to buy 20 units of y, what is the largest number of unts of that she can afford?


1
Expert's answer
2021-02-25T18:17:15-0500
"6z+20y=I"

"z=2y"

"y=\frac {I}{32}"

If z=y


"az+20y=2I"

"ay=\frac {44}{32}I"

"a\frac {I}{32}=\frac{44}{32}I"

"a=44"